A small sphere has a mass of 0.02 kg and the mass of large sphere is 0.4 kg. The center distance between the two sphere is 0.03 m. Calculate the gravitational force on each sphere due to each other.


1
Expert's answer
2021-01-24T20:03:00-0500

The gravitational force on each sphere due to each other can be calculated according to the Newton's law of gravity:


"F = G\dfrac{m_1m_2}{r^2}"

where "m_1 = 0.02kg, m_2 = 0.4kg" are the masses of the spheres, "r = 0.03m" is the distance between spheres, and "G = 6.67\times 10^{-11}N\cdot m^2/kg^2" is the gravitational constant. Thus, obtain:


"F = 6.67\times 10^{-11}\dfrac{0.02\cdot 0.4}{0.03^2} \approx 5.92\times 10^{-10}N"

Answer. "5.92\times 10^{-10}N".